User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Hobart 26m:

Disk VSN:HOB+0015/6000/

Data volume at beginning:0

  • 17:00 UT – started OK (Liza)
  • 23:10 UT – power levels in if3 adjusted. VCs look happier now (Lucia).
  • 05:00 UT – vacuum pressure has been increasing over the last two (03 UT - 05 UT )hours. Now reset the vacuum pump and it is under 3.5 again. Receiver still warm. Possibly bad scans from 202-0303 ongoing due to warm receiver (up to 140K). (Lucia)
  • 11:18 UT – Windstowed (JS)
  • 12:05 UT – Still windstowed (JS)
  • 13:50 UT – Still windstowed (JS)
  • 14:01 UT – Not windstowed! Missed scans 202-1118a to 202-1359a (JS)
  • Drives error at ~16:10. No sources reached between 16:11 and 16:58 (JMc)

Katherine 12m:

Disk VSN: UVLBI-06/2400/

Data volume at beginning: 0

  • 17:00 UT – experiment started OK not without fight with dbbc: the huge formatter-to-GPS clock difference (1E-1) forced to restart and reconfigure DBBC software before observation started (liza)
  • 04:28 UT – fs restarted. Fs has died when extracting a whole lot of data from Monica. Missed scans 202-0408 through 202-0430. 202-0436 should be fine again. Recording was running when fs stopped working; so disk pos now 15 GB ahead. (Lucia)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1697.txt · Last modified: 2015/07/21 16:59 by Jamie McCallum